OpenStudy (ymendoza88):

Find the area between y=4sin(x) and y=4cos(x) over the interval [0, π]

OpenStudy (dumbcow):

cos graph starts above sin graph, then ends below sin graph find point of intersection x* \[A = 4\int\limits_0^{x*} \cos x - \sin x dx + 4 \int\limits_{x*}^\pi \sin x - \cos x dx\]